# Aristotle’s Second-Order Logic

### v11

@@ Conclusion
We have shown how ASOL is the most natural framework to interpret the ’metalogic’ of the Analytics and the logic implicit in the Topics. The structure of the sentence and inferences involved confirms the competency of ancient logicians with regards to quantifier logic and multiple generality. If second-order logic is the most natural framework to read the Organon as a whole then this sheds new light of the whole standard narrative of the development of Western logic suggesting a much greater formal and conceptual continuity between ancient logic and Frege’s system of the Begriffsschrift which is usually equated to second-order logic with a term-forming extension operator. If Frege’s formal system has structural similarities with Aristotelian logic as argued for in this paper, his philosophy of logic and language has been argued (by Bobzien and others ) to have been influenced by the Stoics.
We now discuss in light of the present work the system of argumentation for which the Topics was written. In the Platonic Dialogues (particularly the early262626in a future paper we show the surprising second-order formalizability of the earlier Socratic dialogues as contrasted to the later ones. and middle ones) as well as in Aristotle there is the fundamental supposition that philosophical knowledge can be gained by finding the right definition of terms (under the proviso that the term in question is acknowledged to exist, that is, have non-empty extension.). This concern is not abandoned in the Analytics. The Socratic attitude is that the first step to wisdom consists in testing one’s beliefs which can lead to the valuable cathartic knowledge of the error of one’s beliefs. Thus a major part of philosophical activity involved the testing (i.e. attempt at refutation) of proposed definitions of terms - something which Aristotle acknowledges to be easier than finding correct definitions [^27] A massive amount of arguments were compiled during the eras of the late Academia and first Stoic schools, as testified by accounts of the lives of Chrysippus and Carneades.. In a modern framework how are we to understand (broadly speaking) the dialectical game of the Topics ? In ASOL in which we assume we have a finite collection of constants, functions and predicate symbols (first and second order) which represent the disambiguated vocabulary of a common language. The non-logical axioms we assume are $T$, the topics (or rather a minimal generating set for them ) and $K$ those representing common agreed upon knowledge. The game begins with Alice’s proposal of a definition of a certain term $hAD$. Then Bob basically has to try to derive a contradiction $T+K+hAD\vdash\bot$. The way the game was played apparently involved Alice confirming every move made by Bob, as attested in the Socratic dialogues (where it is rare for the ’opponent’ not to agree with Socrates’ moves). The ’moves’ can be $K$ axioms themselves or inferences, either purely logical or based on a topic. A nice example of this dialectical game in practice is given by Plato’s Lysis (see the ASOL formalization of a key section in appendix 2) although the types of definition considered do not fall into Aristotle’s scheme for monadic predicates terms outlined in book VI of the Topics. As discussed at the end of the previous section, it would be interesting to study the early and middle Platonic dialogues from a second-order perspective as in the present work so as to provide additional evidence for logical competency in dealing with multiple generality among the ancients.
@@
We can of course question if this dialectical game could lead to actual progress in philosophy or to certain or plausible knowledge about reality. In particular the exact scope of $K$ seems to be vague and uncertain; the assumption of the possession of a perfectly disambiguated language is itself implausible (although it should be said that the Topics is acutely aware of this problem). However we should not go to the other extreme of discarding this kind of philosophical practice. Rather what was required is a preliminary radical epistemic, semantic and logical analysis in order to set up a relevant and meaningful system - and such concerns are in no way foreign to the mature Aristotle: things better known to us are not necessarily better known in themselves282828as we saw in section 5, the extra-logic second-order relation of ’being better known’ plays an important role in the Topics: if $A$ is a property of $B$ then it has to be better known than $B$. In book I,1 Aristotle says that a syllogism is a ’proof’ when the premises are ’true and primitive’ or epistemically depending on ’true and primitive’ premises. The adjective ’alethê’ is also translated as ’evident’. Aristotle then clarifies that evident and primitive is what is known through itself and not through another’. All ’scientific principles’ should be self-evident. Other propositions are ’endoksa’, translated sometimes as ’probable’..
@@
As proposed in section 4, in order to gain a greater understanding of the logical, semantic and grammatical content of the Topics we propose the compilation of a list of all the definitions (both accepted and rejected) mentioned in this work and to pay close attention to their grammar and syntax and how this relates to their semantic and logical structure. We also need to clarify in greater detail the concept of predication deployed throughout the work and how it related to the concrete forms of predication given by the predicabilia. Additional indirect arguments for the adequacy of ASOL might be provided by showing that this is also the most natural approach to formalizing Euclid’s Elements, a prototype of which was surely known by Aristotle. We include a short discussion on the first proposition of Euclid in appendix 3.
Appendix 1: some key topics of book IV
@@
Here we list the ASOL interpretation of several important logical topics in book IV. To this end we make use of the list compiled in [16][pp.691-694]. We divide the list into the corresponding chapters of book IV. We introduce an additional second-order predicate $\mathcal{C}AB$ to indicate $A$ and $B$ belong to the same category. We take $\gamma AB$ to encompass also what Aristotle calls ’$A$ participating of $B$’ and ’$A$ receiving the definition of $B$’ (121a). We use extensional containment $A\subset B$ for unspecified ’predication’.
@@
    $\gamma BC\&\sim\gamma AC\rightarrow\sim AB$
@@
    $\gamma BC\&\gamma CA\rightarrow\sim AB$ and also Ch.1,4
@@
    $Ax\&\sim Bx\rightarrow\sim\gamma$ and $Ax\&\gamma BC\&\sim Cx\rightarrow\sim\gamma AB$
@@
9. $\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\mathfrak{g}RS$ (124b)292929if $R$ is one-to-one.
Appendix 2: Plato’s Lysis as an example of ancient Dialectics
We apply ASOL to formalize the dialectics in 212-222 of the Lysis illustrating the naturalness of this approach. As in the Topics there is a second-order function which yields the ’opposite’ of a relation $R$ which we denote by $R^{\bullet}$. For instance if $Lxy$ means ’$x$ loves ’$y$’ then $L^{\bullet}xy$ is to be interpreted as $x$ hates $y$. We suppose with Plato that $R^{\bullet\bullet}=R$ and that $R=S\rightarrow R^{\bullet}=S^{\bullet}$. Also that $Rxy\rightarrow\sim R^{\bullet}xy$. We assume
$(Sxy\leftrightarrow Txy)\leftrightarrow(S^{\bullet}xy\leftrightarrow T^{\bullet}xy)$
We use a Zalta-type approach to Platonic forms whereby we gather one or more predicates into an abstract first-order object. Thus for instance $[X:X=\Phi]$ would be the singleton $[\Phi]$ which represents the abstract object encoding only $\Phi$ (see [25] for more information about the concept of encoding considered here).
We fix a subset $\mathcal{C}$ of all sentences of ASOL which we call sentences of common assent together with a set $\mathcal{T}$ of sentences considered to be true.
The Lysis concerns the definition of the binary relation $\Phi xy$ ($y$ is $x$’s friend). We are given a relation $Lxy$ ($x$ loves $y$). We now attempt to formalize the essential structure of the dialogue in 212-222.
Socrates begins by inquiring concerning the logical relationship between $\Phi$ and $L$. Lysis assents to the first hypothesis for a definition
$1.\enspace Lxy\enspace\&\enspace Lyx\leftrightarrow\Phi xy\quad\text{Hyp}$
Socrates then makes the deduction
$2.\enspace\sim Lyx\rightarrow\sim\Phi xy\quad\text{1}$
and brings out several sentences of $\mathcal{C}$ such as
$3.\enspace\exists x,y\,Man\,x\enspace\&\enspace Horse\,y\enspace\&\sim Lyx\enspace\&\enspace Lxy\enspace\&\enspace\Phi xy$
and $\exists x.Man\,x\enspace\&\enspace\Phi x\,wisdom\enspace\&\sim L\,wisdom\,x$ which refute 1 and thus we arrive at the conclusion
[…]
Then Socrates puts forward another hypothesis:
$5.\enspace Lxy\leftrightarrow\Phi xy\quad\text{Hyp}$
Applying (Op) we get
$6.\enspace L^{\bullet}xy\leftrightarrow\Phi^{\bullet}xy$
But clearly $\exists x,yLxy\enspace\&\enspace L^{\bullet}yx$. Hence we deduce that
$7.\enspace\exists x,y\,\Phi xy\enspace\&\enspace\Phi^{\bullet}yx$
There will be people who are friends of their enemies and enemies of their friends. The negation of 7 is in $\mathcal{C}$ and so hypothesis 5 is also rejected and we have
[…]
Socrates then makes an appeal to the wisdom of the poets and brings forth the predicates $Hxy$(likeness) and $Kx$ (good) alongside their opposites. For $H$ opposition and negation seem to coincide.
Likeness can be extended to $\mathfrak{H}xyz$ which is to be interpreted as ’likeness of $x$ to $y$ according to $z$’ and thereby resolve contradictions. So too for other relations. The following it put forth
$9.\enspace Hxy\leftrightarrow\Phi xy\quad\text{Hyp}$
Then follows interesting propositions involving $H$:
$10.\enspace\sim Hxx\rightarrow\forall y\sim Hxy$
$11.\enspace K^{\bullet}x\rightarrow\sim Hxx\enspace\&\enspace Kx\rightarrow Hxx$
From these propositions (assumed to be part of $\mathcal{C}$ or $\mathcal{T}$) Socrates then deduces (assuming $Kx\enspace\&\enspace Ky\rightarrow Hxy$) that
$12.\enspace Kx\enspace\&\enspace Ky\rightarrow\Phi xy\enspace\&\enspace K^{\bullet}x\enspace\&\enspace K^{\bullet}y\rightarrow\sim\Phi xy$
So bad people cannot friends.
Let $Pxyz$ mean $x$ can do $z$ to $y$. Socrates assumes
$13.\enspace Hxy\rightarrow\forall z(Pxyz\rightarrow Pxxz)$
Let $\mathfrak{K}xy$ be defined as $\exists z\sim Pxxz\enspace\&\enspace Pyxz$. The axiom of the self-sufficiency of the good is:
$Kx\rightarrow\sim\exists y\mathfrak{K}xy$
Socrates accepts the pragmatic condition for friendship
$14.\enspace\Phi xy\rightarrow\mathfrak{K}xy$
from which we deduce
$15.\enspace\Phi xy\rightarrow\sim Hxy$
and hence hypothesis 9 is refuted
[…]
Next, inspired by Hesiod, Socrates proposes:
$17.\enspace\sim Hxy\leftrightarrow\Phi xy$
Socrates then draws the obvious conclusion from 17:
$18.\enspace\Phi[\Phi][\Phi^{\bullet}]\enspace\&\enspace\Phi[\Phi^{\bullet}][\Phi]$
Friendship is friends with enmity and likewise justice is friends with injustice and so forth. Since this must be rejected we get
[…]
There can be a middle ground or mixture between opposites or a least a ground which belongs to neither of the opposites. Socrates proposes things which are neither good nor evil $\sim Kx\enspace\&\sim K^{\bullet}x$.
We denote this predicate by $Mx$. There are thus three classes of beings
$20.\enspace\forall x(Kx\vee K^{\bullet}x\vee Mx)$
We can use these classes to propose a definition of the form
$\mathbb{P}x\enspace\&\enspace\mathbb{Q}x\rightarrow\Phi xy$
where $\mathbb{P}$ and $\mathbb{Q}$ are among $K,K^{\bullet}$ and $M$. Since $x$ such that $K^{\bullet}x$ is ruled out from entering into $\Phi$ it follows from 15 and (K) that the only possibility is
$21.\enspace Mx\enspace\&\enspace Ky\rightarrow\Phi xy\quad\text{Hyp}$
Let $X$ be a $n$-ary predicate variable and $Y$ a $n+1$-ary predicate variable. We introduce the second-order relations $X\prec_{n}Y$ which must satisfy the axiom
$X\prec_{n}Y\rightarrow Xx_{1}...x_{n}\leftrightarrow\exists yQx_{1}...x_{n}y$
We consider $\Phi\prec_{2}\Phi^{+}$ in which $\Phi^{+}xyz$ is interpreted as $y$ is friend to $x$ for the sake of $z$. Socrates states that
$22.\enspace\Phi^{+}xyz\rightarrow\Phi xz$
and that the primary instance of $\Phi xy$ we are interested in is that in which $\Phi^{+}xyy$.
Apparently Socrates rules out an endless sequence $\Phi^{+}xya_{1},\Phi^{+}x,a_{1},a_{2},\Phi^{+}xa_{2}a_{3},...$ as well as a cycle such as $\Phi^{+}xyz,\Phi^{+}xzy$. Let $\mathfrak{F}xy$ be by definition $\Phi^{+}xyx$.
We also define $\Phi\prec_{2}\Phi^{a}$ where $\Phi^{a}xyz$ expresses that $z$ is a cause for $\Phi xy$ in the way that sickness is the cause of a doctor being a friend to a man. Socrates states that
$23.\enspace\Phi^{a}xyz\rightarrow K^{\bullet}z$
But
$24.\enspace\exists x,y,z\sim K^{\bullet}z\enspace\&\enspace\Phi^{a}xyz$
so 23 must be rejected.
There is the predicate $Bxy$, $y$ belongs to $x$. Socrates proposes
$25.\enspace Bxy\rightarrow\Phi xy$
but
$26.\enspace Kx\rightarrow(Bxy\rightarrow Ky)$
$27.\enspace Mx\rightarrow(Bxy\rightarrow My)$
$28.\enspace K^{\bullet}x\rightarrow(Bxy\rightarrow K^{\bullet}y)$
Hence since there is a $y$ such that $Bxy$ and $K^{\bullet}x$ then by $24$ we will have $\Phi xy$ which contradicts 12. Hence we reject 25.
Appendix 3: the first proposition of Euclid
Here we propose how a minimal fragment of Euclid (enough to formulate and prove the first proposition of the first book) might be formalized in ASOL.
Individuals correspond to points in the plane. There are no magnitudes as primitive sorts. Our primitives are $\Lambda Lxy$ for $L$ a monadic predicate variable, $\Sigma CxL$ for $C$ and $L$ monadic predicate variables and $A<B$ for $A$ and $B$ monadic predicate variables. The intended meaning of $\Lambda$ is: $L$ is a line with extremities $x$ and $y$, of $\Sigma$: $C$ is a circle with center $x$ and of radius equal to the length of line $L$ and $A<B$: line $A$ is shorter than line $B$. We employ, alongside standard equality for individuals, an extensional second-order equality on monadic predicate terms defined as
$A=B\leftrightarrow\forall x(Ax\leftrightarrow Bx)$
We define $A\equiv B\leftrightarrow A\nless B\&B\nless A$. This is to be interpreted as $A$ and $B$ having the same length (when lines) and should not be confused with extensional equality.
We assume that we can form ’intersections’ and ’unions’ and ’complements’ of monadic predicate terms in the expected way using the $\lambda$. For instance
$A\cup B=\lambda x(Ax\vee Bx)$
We can now define a triangle:
$\Delta Fabc\leftrightarrow a\neq b\neq c\&(\forall L_{1}L_{2}L_{3}(\Lambda L_{1}ab\&\Lambda L_{2}bc\&\Lambda L_{3}ca\rightarrow F=L_{1}\cup L_{2}\cup L_{3})$
and equilateral triangle
$\Delta^{E}Fabc\leftrightarrow\Delta Fabc\&(\forall L_{1}L_{2}L_{3}(\Lambda L_{1}ab\&\Lambda L_{2}bc\&\Lambda L_{3}ca\rightarrow L_{1}\equiv L_{2}\equiv L_{3})$
We assume the basic axioms
$\forall xy\exists L\Lambda Lxy$
$\forall xR\exists C\Sigma CxR$
$\forall xRCy\,\Sigma CxR\&Cy\rightarrow(\forall L\,\Lambda Lyx\rightarrow L\equiv R)$
and the axiom which Euclid took apparently for granted:
$\forall xyC_{1}C_{2}R(\Sigma C_{1}xR\&\Sigma C_{2}yR\&\Lambda Rxy\rightarrow\exists zC_{1}z\&C_{2}z)$
We now can prove
Theorem 6.1
$\forall xyL\,\Lambda Lxy\rightarrow\exists zT\,\Delta^{E}Txzy$
The proof in ASOL is surely shorter and more natural, closer to Euclid’s own proof, than a more convoluted first-order approach.
Appendix 4: Presentation of ASOL and ESOL
ASOL is a special case of extended second-order logic (ESOL) defined as follows. We are given a collection of individual variables $x,y,z,...$ and individual constants $a,b,c,...$ together with a collection of $n$-ary predicate term variables $X^{(n)},Y^{(n)},Z^{(n)},...$ (we omit the superscript if the context is clear) for $n\geq 0$. Furthermore we are given a collection of predicate term symbols $\alpha^{(\tau)},\beta^{(\tau)},\gamma^{(\tau)},...$ where $\tau$ consists of a (possibly empty) list $(n_{1},...,n_{m})$ of integers $n_{i}\geq-1$ called a signature, where we likewise omit the superscript if the context is clear. It is also convenient to consider that we have a collection of function symbols $f^{(\tau,n)},g^{(\tau,n)},...$ where $\tau$ is a signature and $n\geq-1$. For $n=-1$ these are called first-order functional symbols. Otherwise they are called second-order function symbols.
If $\alpha^{(\tau)}$ is a predicate term symbol with $\tau=(n_{1},...,n_{m})$ non-empty then $\alpha\mathfrak{t}_{1}....\mathfrak{t}_{m}$ is a formula where if $n_{i}\geq 0$ then $\mathfrak{t}_{i}$ is a $n_{i}$-predicate term and if $n_{i}=-1$ then $\mathfrak{t}_{i}$ is an individual term. If $\tau$ is empty then $\alpha$ is considered a 00-predicate term.
If $\Phi$ is a formula and $x_{1},...,x_{n}$ is a list of individual variables (possible empty) then $\lambda x_{1}...x_{n}\Phi$ is a $n$-predicate term (for the empty list we put $n=0$ and write $\lambda\Phi$.
If $X^{(n)}$ is a predicate term variable then it is a $n$-predicate term. If $\alpha^{(\tau)}$ is a predicate term symbol such that $\tau=(k_{1},...,k_{n})$ with $k_{i}=-1$ then $\alpha^{(\tau)}$ is a $n$-predicate term303030we can also call $\alpha$ in this case a $n$-predicate constant.
An individual constant or variable is an individual term.
If $f^{(\tau,n)}$ is a function symbol then $f\mathfrak{t}_{1}...\mathfrak{t}_{m}$ is an individual term (for $n=-1$) or $n$-predicate term (for $n\geq 0$) where $\mathfrak{t}_{1}...\mathfrak{t}_{m}$ satisfy the condition relative to $\tau$ as above.
The natural deduction rules consist of the standard rules for the first-order classical predicate calculus:
The rules are as follows:
[…]
the four first-order quantifier rules:
[…]
We have the analogous second-order quantifier rules
[…]
with analogous provisos.
Finally there are introduction and elimination rule for $\lambda$-predicate terms for which the reader may consult [15].
In Prawitz[15] versions 1 and 2 are proof-theoretically equivalent (for formulas without $\lambda$), the difference being that in version 1 the rules $\lambda I$ and $\lambda E$ are incorporated into second-order quantifier rules.
Why do we consider ESOL as being an extension of second-order logic rather than a fragment of third-order logic ? We are of the view that what defines the ’order’ of a logic is what is quantified over and what kind of corresponding variable there is. First-order logic quantifies over individuals and has individual variables. Second-order logic quantifiers over entities corresponding to $n$-ary predicates with individual arguments and has the corresponding $n$-ary predicate variables. Third-order logic thus would have to quantify over entities corresponding to predicates which can involve both individual arguments and $n$-ary predicates of individuals arguments and furthermore it would require variables to represent these last. Such a quantification and such variables are lacking in ESOL although the predicate term symbols $\alpha^{(\tau)}$ can be considered in some sense as being ’third-order’ when involving $n$-predicate term arguments.
There are some extensions and modifications of ESOL which can make it even more adequate to formalize Aristotle’s reasoning. These also are interesting in their own right. Thus our project will involve the following:
  1. 1.
[…]
    The reason for this investigation is to develop my alternative proposal for an interpretation of Aristotle’s demonstrative technique of ekthesis which calls for a generalized indefinite version of the Peano iota operator (while the Peano operator corresponds to a definite article this construction corresponds to the indefinite article). If we have that there exists an A, then we can form the term ’an A’ so as to be able to proceed thus: let a be an A. An example of a natural language expression of the rule above: every man has a father, hence, every man’s father is the father of that man (or: if every man has a father then a father of a man is that man’s father).
  2. 2.
    In natural language discourse most (if not all) quantifiers appear to be bounded. This is true of Aristotle’s reasoning and this approach has proven itself useful in contemporary applications of dependent type theory. Thus we will investigate alternative forms of the natural deduction rules for ESOL in which all quantifiers are bounded $\forall x(\Phi(x)\rightarrow\Psi(x))$ (we write this $\forall_{x\Phi}\Psi$). In this alternative natural deduction system Modus Ponens is incorporated into the quantifier elimination rules, much as the $\lambda$ rules are incorporated into the quantifier rules in Prawitz’s first version of second-order logic. Some of the rules might look like:
[…]
    Rule $\forall^{1}E$ is of course the famous syllogism involving singular propositions (not in fact found in Aristotle’s Organon): all men are mortal, Socrates is mortal, therefore Socrates is mortal. To bring the logic closer to natural language we could also construct an alternative system in which all formulas (including the formulas appearing in the rules) are closed formulas. This would require for instance Modus Ponens being incorporated into the new quantifier laws or vice-versa. For instance we would have to replace MP we rules like:
[…]
    If we combine this with bounded quantifiers then it seems we would require instead of MP extensions like:
[…]
    which is just the old Barbara syllogism.
  3. 3.
    Recall that starting from $\forall^{2},\forall^{1},\rightarrow$ we can define all the other connectives. We can define ESOLat in analogy to atomic polymorphism for system $F$ by restricting $\forall E$ to predicate terms without $\lambda$. We can embed system $F$ ($\lambda 2$) into ESOL. Furthermore this embedding can be used to show that ESOL is undecidable due to $F$ being undecidable. Does the undecidability of $F_{at}$[1] carry over to ESOLat ?
  4. 4.
    There is a certain redundancy in our definition of formulas and terms in ESOL. In the definition of formula in ESOL the construction of a formula via the application of an $n$-predicate term to a list of individual variables $Xt_{1}...t_{n}$ can be considered as a special case of the application of predicate symbols (see the previous definition) via a special predicate term symbol $\Delta^{n}$ with signature $(n,-1,...,-1)$ satisfying axiom:
    $Xt_{1}....t_{n}\leftrightarrow\Delta^{n}Xt_{1}...t_{n}$
    We propose to explore this alternative presentation of ESOL. Another topic is to clarify the relative strength and relationship between ESOL and multi-sorted first-order logic as well as Bealer’s intensional logic (such as presented, for instance, in Quality and Concept[2]).
On a philosophical side we plan to develop the following arguments for the philosophical interest as well as applicability of ESOL.
  1. 1.
    One reason to consider ESOL as a philosophical logic arises when we are confronted with situations in which we wish to distinguish intention from extension, the meaning of a concept from its extension. For example for Aristotle the concept ’able to laugh’ and ’rational animal’ both have humans as their extension but their meanings are clearly different. It is necessary to have an equality predicate defined not only for individuals but for $n$-ary predicate terms, in particular monadic predicate term constants. This equality $C^{1}=D^{1}$ need not be equivalent to $\forall x(Cx\leftrightarrow Dx)$.
  2. 2.
    Zalta’s Object Logic [25], which in one presentation is a second-order logic with fusions of concepts into abstract objects, provides an alternative interpretation of Fregean Arithmetic. It is based on a primitive encoding relation $enc\,tT^{1}$ which involves an individual term $t$ and 1-ary predicate term $T^{1}$. Object Logic is thus a particular case of ESOL.
  3. 3.
    In order to formalize propositional attitudes it is convenient to have predicates which can take 00-ary predicate terms as arguments. ESOL meets this need. Note that in ESOL a 00-ary predicate term $T^{0}$, for instance $\lambda\Phi$ for some sentence $\Phi$ can be construed as expressing the meaning of $\Phi$ while $T^{0}()$ expresses the judgment involving a truth-value claim. Thus the relation $K$ of agent $a$ knowing sentence $\Phi$ would be expressed as $Kx\lambda(\lambda\Phi)())$.
  4. 4.
    In certain situations (as in ’Zig-Zag and Fregean Arithmetic’[8]) we want to distinguish between predicative and non-predicative predicates, concepts which are ’too complicated’ to have extensions (analogously to how some extensions are ’too complicated’ to be the extension of a single concept). This again calls for a monadic predicate having predicate term arguments, for instance $\pi X^{(1)}$ expressing that $X^{(1)}$ is predicative.
  5. 5.
    We can also consider functions taking predicate terms as arguments and forming either individual terms or predicate terms. The later case could be used to interpret first and second-order multi-modal logic as already suggested in this paper.

### v17

@@ Conclusion
We have discussed in some detail how ESOL might be used to formalize the Topics. If ESOL is the most natural framework to read the Organon as a whole then this sheds new light on the development of Western logic, suggesting a much greater formal and conceptual continuity between ancient logic and Frege’s system of the Begriffsschrift, usually equated to second-order logic with a term-forming extension operator.
@@
Regarding section 6 we can of course question if this dialectical game could lead to actual progress in philosophy or to certain or plausible knowledge about reality. In particular the exact scope of $K$ seems to be vague and uncertain; the assumption of the possession of a perfectly disambiguated language is itself implausible (although it should be said that the Topics is acutely aware of this problem). However we should not go to the other extreme of discarding this kind of philosophical practice. Rather what was required is a preliminary epistemic, semantic and logical analysis of our concepts and language - and such concerns are in no way foreign to the mature Aristotle: things better known to us are not necessarily better known in themselves232323as we saw in section 6, the extra-logic second-order relation of ’being better known’ plays an important role in the Topics: if $A$ is a property of $B$ then it has to be better known than $B$. In book I,1 Aristotle says that a syllogism is a ’proof’ when the premises are ’true and primitive’ or epistemically depending on ’true and primitive’ premises. The adjective ’alethê’ is also translated as ’evident’. Aristotle then clarifies that evident and primitive is what is known through itself and not through another’. All ’scientific principles’ should be self-evident. Other propositions are ’endoksa’, translated sometimes as ’probable’.. The theory of the Categories is already assumed in the Topics. We can ask if Aristotle envisioned the project of a systematic and complete classification of all terms (the entire semantic universe) according to the theory of definition discussed in this paper, starting from the supreme genera. Such a project (which contrasts in certain aspects with that of Leibniz’ characteristica) bears some similarity to the goal of Roget’s famous Thesaurus or of modern lexicological projects.
@@
As proposed in section 4.3, in order to gain a greater understanding of the logical, semantic and grammatical content of the Topics it would be useful to compile a list of all the definitions (both accepted and rejected) mentioned in this work and to pay close attention to their grammar and syntax and how this relates to their semantic and logical structure. We also need to clarify in greater detail the concept of predication deployed throughout the work and how it related to the concrete forms of predication given by the predicabilia.
We section 8 have argued for and provided evidence that Aristotle was in possession of a theoretical logic completely adequate for reasoning about quantified sentenced containing multiple generality (together with the basic rules for implication and reductio ad absurdum). This thesis was further corroborated with material from Galen and Boethius. This work can be seen to parallel in the spirit of its historical approach to the work of Bobzien[5] and Bobzien and Shogry[6] on the sophistication of the deductive system of Stoic propositional logic and the adequacy of this logic to express multiple generality. While much remains to be studied and explored regarding quantifier reasoning and multiple generality in medieval logic the conclusion which seems to suggest itself is that our standard view of the history of western logic needs to be somewhat revised so that a greater continuity is acknowledged between ancient logic and the development of quantifier logic by Frege and Peirce as well as the calculus of relations of Schröder. A remarkable consequence is that if our interpretation and reconstruction of Boethius’ logic given in this paper be correct then it furnishes some vital clues to better understand Kant’s conception of logic as expounded in the Critique of Pure Reason and frees Kant to a certain extent (in so far as he would have been acquainted with Stoic logic and the writings of Boethius) from logical naivety or the imprisonment in first-order monadic logic, a logic unsuitable for mathematics and physics as well as philosophy.
Appendix 1: key topics of book IV
@@
Here we list the ESOL interpretations of several important logical topics in book IV. To this end we make use of the list compiled by Owen in [15][pp.691-694]. We divide the list into the corresponding chapters of book IV. We introduce an additional predicate $\mathcal{C}AB$ to indicate $A$ and $B$ belong to the same category. We take $\gamma AB$ to encompass (as already mentioned) what Aristotle calls ’$A$ participating of $B$’ and ’$A$ receiving the definition of $B$’ (121a). We use extensional containment $A\subset B$ for unspecified ’predication’.
@@
    $\gamma BC\&\sim\gamma AC\rightarrow\sim\gamma AB$
@@
    $\gamma BC\&\gamma CA\rightarrow\sim\gamma AB$ and also Ch.1,4
@@
    $Ax\&\sim Bx\rightarrow\sim\gamma AB$ and $Ax\&\gamma BC\&\sim Cx\rightarrow\sim\gamma AB$
@@
9. $\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\gamma_{2}RS$ (124b)242424if $R$ is one-to-one.